OpenStudy (anonymous):

\[\int\limits_{0}^4\int\limits_{\sqrt{y}}^{2}{\sqrt{x^{3}+1}} dxdy\]

OpenStudy (blockcolder):

Is that lower bound on x \(\Large \sqrt{3}\)?

OpenStudy (anonymous):

it's on \[\sqrt{y}\]

OpenStudy (blockcolder):

Right. Our first step here is to see that we really can't evaluate \(\large\int \sqrt{1+x^3} dx\) so we can't evaluate the double integral as it is. One possible remedy that usually works is to reverse the order of integration. But first, give me a sketch of the region we're integrating over.

OpenStudy (anonymous):

∫1+x3−−−−−√dx

OpenStudy (anonymous):

they don't give me a sketch...they just say evaluate the integral...i tried to do it as they gave me but once i integrate with respect to x it gets messy... it becomes \[2/9\int\limits_{0}^{4}27/4-((y \sqrt{y}+1)^3/2)/y dy\]

OpenStudy (anonymous):

it's supposed to be raised to the 3/2 and not raised to 3 and then divided by 2

OpenStudy (blockcolder):

Yeah, that's because you can't directly evaluate \(\Large \int \sqrt{1+x^3}\ dx\), as I said. However, you CAN evaluate \(\Large \int \sqrt{1+x^3} dy\), and that's what we're gonna do. If they don't provide a sketch, it's up to you to infer what the region is based on the bounds of the integrals. For now, here's the sketch: